Help needed!

Kathi and Robert Hawn had a pottery stand at the annual skipcack craft fair. They sold some of their pottery at the original price of $11.50 for each piece, but later decreased the price by $2. If they sold all 82 pieces and took in $821.00, find how many pieces they sold at the original price and how many pieces they sold at the reduced price?
The number sold at $11.50 is___
The number sold at $9.50 is ____

Answer by rothauserc(4718)

let x be the number they sold at $11.50 and y be the number they sold at $9.50, then we have
x + y = 82
x*11.50 + y*9.50 = 821.00
solve the first equation for x
x = 82-y
substitute for x in second equation
11.50*(82-y) + y*9.50 = 821.00
943 - 11.50y +9.50y = 821.00
-2.00 = -122.00
y = 61
x = 82 - 61 = 21
The number sold at $11.50 is 21
The number sold at $9.50 is 61
check answers
21*11.50 + 61*9.50 = 821.00
241.50 + 579.50 = 821.00
821.00 = 821.00
our answer checks :-)